OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2+3

OpenStudy (anonymous):

Well.....it is x^2 so you have 4^2 that equals 16.

OpenStudy (anonymous):

So, 16 + 3 = 19.

jimthompson5910 (jim_thompson5910):

x^2 means x*x So 4^2 = 4*4 = 16 and NOT 8

OpenStudy (anonymous):

Oh! >.< Gah!, Thanks!